# Khushbu Pahwa3, Jian Ding1, Mohamed Elhoseiny1

### v4

@@ Abstract
Short TGIF-QA (Jang et al., 2017) 165.2 K 72.0 K 0.05 60.00
MSRVTT-QA (Xu et al., 2017) 243.6 K 10.0 K 0.25 416.6
MV-Bench (Li et al., 2024b) 4.0 K 3.6 K 0.27 16.38
TVQA (Lei et al., 2019) 152.5 K 21.8K 1.27 461.20
Long Activity-QA (Yu et al., 2019) 58.0 K 5800 3.00 290.00
Egoschema (Mangalam et al., 2023) 5.0 K 5063 3.00 253.15
LongVideoBench (Wu et al., 2024) 6.7 K 3763 7.88 494.21
Moviechat (Song et al., 2023) 14.0 K 1000 9.40 156.67
MLVU (Zhou et al., 2024) 3.1 K 1730 15.50 446.92
MoVQA (Zhang et al., 2023b) 21.9 K 100 16.53 27.55
Video-MME (Fu et al., 2024) 2.7 K 900 16.97 254.55
Very Long LVBench (Wang et al., 2024b) 1.6 K 103 68.35 117.33
@@ 1 Introduction
Recent progress in Multimodal large language models (MLMMs) has enabled impressive performance on image and short-video understanding by jointly processing visual and textual inputs (Ataallah et al., 2024a; Zhu et al., 2023; Zhang et al., 2023a; Chen et al., 2023; Lin et al., 2023; Liu et al., 2023; Maaz et al., 2023). To push toward richer video comprehension, several models have extended the context length of underlying LLMs, enabling them to process longer videos  (Bai et al., 2023, 2025; Li et al., 2024a). However, current public benchmarks remain limited in both scope and realism. Most are restricted to short clips (a few seconds to minutes), and even larger-scale efforts like LVBench, with one-hour videos, lack narrative complexity and modality diversity (Wang et al., 2024b). In real-world scenarios, long-form videos especially movies and TV shows present fundamentally different challenges. These formats feature rich, structured storytelling with long-range temporal dependencies, evolving character arcs, and intricate causal relationships (Zhang et al., 2023b; Lei et al., 2019; Song et al., 2023) . Unlike short videos, they demand high-level skills such as intent inference, multi-event linking, and thematic summarization. Humans can easily integrate dispersed cues across modalities and time; current MLMMs cannot.
@@ 2 Related Work
Recently, many multimodal large language models have aimed to extend their context lengths, such as Qwen2.5-VL (Bai et al., 2025), Qwen2-VL (Wang et al., 2024a), and LongVU (Wu et al., 2024), which can process videos over an hour long. This highlights the growing need for effective benchmarks. In contrast, short video benchmarks have been widely explored, covering tasks like moment retrieval (Liang et al., 2024; Lei et al., 2021), action classification (Caba Heilbron et al., 2015; Carreira and Zisserman, 2017; Soomro et al., 2012), video reasoning (Xie et al., 2024; Xiao et al., 2021), and video QA (Xu et al., 2017; Lei et al., 2019; Jang et al., 2017). Among longer video benchmarks, TVQA (Lei et al., 2019) is an early example with over 152.5 K QA pairs across 21.8 K clips (1.27 min avg.), totaling 461.2 hours. More recent datasets like MovieChat-1K(Song et al., 2023) offer 1,000 movie clips (9.4 min avg.) with 14K annotations. CRAFT(Ates et al., 2020) and GazeVQA (Ilaslan et al., 2023) extend to causal and embodied reasoning. Other long-form efforts include MLVU (Zhou et al., 2024), MoVQA (Zhang et al., 2023b), and Video-MME (Fu et al., 2024), though their average durations peak at 17 minutes. LVBench(Wang et al., 2024b) stands out with 1-hour videos, but includes only 103 clips totaling 117 hours. Domain-specific datasets like Ego4D QA on EgoSchema(Mangalam et al., 2023) focus on first-person views but use short (3-minute) clips. Motivated by these gaps, our benchmark evaluates grounding (temporal/visual anchoring) and reasoning (causality, narrative) skills to reflect human-like video understanding. As shown in Table 1, our dataset supports near-hour-long videos and offers the most significant total duration ($\sim$1K hours) and the largest QA-pair count in terms of long video understanding (87.7K).
@@ 3.2.2 Generating Question-Answer Pairs
Why Transcripts? Transcripts offer richer context than subtitles, including scene descriptions and character actions (e.g., “Monica and Chandler enter the café”). These cues allow for generating deeper, temporally grounded questions that reflect both visual and structural aspects of the story. For more details, see B.4 in the supplementary. For the TV shows, transcripts were sourced from reputable platforms such as Forever Dreaming and manually verified and aligned with the video. For movies, we utilize the transcripts provided by MovieNet Huang et al. (2020), which have been pre-processed to ensure accurate alignment with both dialogue and subtitles, as detailed in Section A.5 Huang et al. (2020).
@@
Character Tracking: To track character actions in the video, we input the transcript and the video summary into GPT-4o, which generates question-answer pairs focused on identifying the sequence of actions performed by the main characters. Deep Context Understanding: For the deep context understanding question-answer pairs, we employed GPT-4o with the transcript and the summary to generate questions about the hidden motivations and social cues. Global Appearance: Our pipeline, as shown in Figure 2, integrates person detection Khanam and Hussain (2024) and face-matching using InsightFace Guo et al. (2021) to identify characters in video frames. Reference images of the main cast were collected from IMDb and matched with detected faces, using an empirically determined threshold to filter out unmatched individuals. Cropped character images were saved chronologically in timestamped folders. To identify unique outfits, we extracted visual features using DINO-v2 Oquab et al. (2024) and removed redundant frames based on similarity scores, reducing approximately 200 frames to 20 while conservatively preserving visual diversity. The remaining images were then processed with GPT-4o OpenAI (2024) to generate outfit descriptions. Multiple-choice questions were constructed by varying the sequence of outfits. In cases where a character wore a single outfit throughout, distractor options featuring incorrect outfits were introduced to maintain question complexity. Summarization & Spoiler Questions: We curate expert-written IMDb IMDB summaries for human-level summarization. Spoiler questions are sourced from flagged plot discussions, focusing on twists or endings.
@@ 3.2.3 Data Integrity and Diversity
1) Redundancy Removal. To avoid question duplication, we compute pairwise textual similarity using BGE-M3 embeddings Chen et al. (2024a) and remove highly similar questions (see Section B.5 for details). Additionally, since movies often contain over 100 events, many of which are redundant, we apply the same similarity-based filtering to retain only 20 diverse and representative events per movie.
@@ 3.3 Data Statistics
InfiniBench ultimately comprises 87.7K questions over videos averaging 53 minutes, with some running as long as 201 minutes, making it the largest and most challenging benchmark for long-video understanding. InfiniBench includes 923 episodes from six TV shows (Lei et al., 2019) based on the long-form version in GoldFish Ataallah et al. (2024b) and filtered 296 movies that have subtitles from MovieNet (Huang et al., 2020). These movies offer a broad and challenging foundation for benchmarking long-video understanding. Figures 4 and 4 illustrate the per-skill distribution, source breakdown (TVQA (Lei et al., 2019), MovieNet (Huang et al., 2020)), and duration statistics across the dataset. We humanly verified 20% of the data, ensuring fair coverage across TV shows and movies. The remaining 80 % is available for training video models. The verified portion is further divided into test and validation splits, 10% each. All reported results and ablation studies are conducted using the test split.
@@ 4.1 Models
Open-Source Models We benchmark eleven leading open-source MLLMs, each evaluated using their official implementations and following their official best practices and configuration, e.g., recommending sampling strategy and frame-rate. These models vary in architecture, context length, and vision-language alignment strategies, offering a diverse perspective on the current research landscape: Qwen2.5-VL Bai et al. (2025), Qwen2-VL Wang et al. (2024a), InternVL3 Zhu et al. (2025b), InternVL2.5 Chen et al. (2024b), InternVL2InternVL (2), LLaVA-OneVision (Li et al., 2024a), LongVU (Shen et al., 2024),VideoChat-Flash Li et al. (2025), InternLM-XComposer-2.5 Zhang et al. (2024), Goldfish Ataallah et al. (2024b), and MiniGPT4-videoAtaallah et al. (2024a)
Commercial Models We also evaluate two cutting-edge proprietary models via their official APIs: GPT-4o OpenAI (2024) and Gemini 2.0 Flash  Gemini (2024). These models represent the frontier of closed-access MLLMs with advanced capabilities in video, text, and multi-turn interaction. Their inclusion offers a valuable benchmark ceiling for current open-source efforts.
@@ 4.4.1 Self-enhancement Bias
The main results in Table 3 use GPT-4o as the judge model, which raises the potential concern of self-enhancement bias Zheng et al. (2023), since GPT-4o serves as both the task model and evaluator. However, we argue that such bias is unlikely in our setup due to the clear separation between generation and evaluation. Specifically, during inference, the model accesses both video and subtitles to generate answers, whereas the judge model evaluates responses using only the generated text and reference answers. This separation in goals, modalities, and inputs minimizes the risk of any model-specific advantage transferring across tasks.
@@ 6 Limitations
Despite its strengths, our question–answer generation pipeline currently depends on the availability of transcripts, limiting its use to movies and TV shows. To broaden applicability, we will develop a more general pipeline that can handle arbitrary videos without preexisting transcripts. We also plan to leverage our dense, skill‐based annotations to create multi‐video benchmarks, such as season‐level evaluations, that extend each skill across related clips. Furthermore, our training split can support retrieval‐based MLLM methods that focus on relevant frames (inspired by Video-XL (Liu et al., 2025)), and we aim to explore long-video reasoning approaches, such as those based on the GRPO algorithm Shao et al. (2024), using samples from our dataset.

### v5

@@ Abstract
Short TGIF-QA (jang2017tgifqa) 165.2 K 72.0 K 0.05 60.00
MSRVTT-QA (xu2017video) 243.6 K 10.0 K 0.25 416.6
MV-Bench (li2024mvbench) 4.0 K 3.6 K 0.27 16.38
TVQA (lei2019tvqa) 152.5 K 21.8K 1.27 461.20
Long Activity-QA (yu2019activitynetqa) 58.0 K 5800 3.00 290.00
Egoschema (mangalam2023egoschema) 5.0 K 5063 3.00 253.15
LongVideoBench (wu2024longvideobench) 6.7 K 3763 7.88 494.21
Moviechat (song2023moviechat) 14.0 K 1000 9.40 156.67
MLVU (zhou2024mlvucomprehensivebenchmarkmultitask) 3.1 K 1730 15.50 446.92
MoVQA (zhang2023movqa) 21.9 K 100 16.53 27.55
Video-MME (videomme) 2.7 K 900 16.97 254.55
Very Long LVBench (wang2024lvbench) 1.6 K 103 68.35 117.33
@@ 1 Introduction
Recent progress in Multimodal large language models (MLMMs) has enabled impressive performance on image and short-video understanding by jointly processing visual and textual inputs (minigpt4_video; zhu2023minigpt; video_llama; minigptv2; video-llava; llava; maaz2023video). To push toward richer video comprehension, several models have extended the context length of underlying LLMs, enabling them to process longer videos  (bai2023qwenvl; bai2025qwen25vltechnicalreport; li2024llavaonevisioneasyvisualtask). However, current public benchmarks remain limited in both scope and realism. Most are restricted to short clips (a few seconds to minutes), and even larger-scale efforts like LVBench, with one-hour videos, lack narrative complexity and modality diversity (wang2024lvbench). In real-world scenarios, long-form videos especially movies and TV shows present fundamentally different challenges. These formats feature rich, structured storytelling with long-range temporal dependencies, evolving character arcs, and intricate causal relationships (zhang2023movqa; lei2019tvqa; song2023moviechat) . Unlike short videos, they demand high-level skills such as intent inference, multi-event linking, and thematic summarization. Humans can easily integrate dispersed cues across modalities and time; current MLMMs cannot.
@@ 2 Related Work
Recently, many multimodal large language models have aimed to extend their context lengths, such as Qwen2.5-VL (bai2025qwen25vltechnicalreport), Qwen2-VL (wang2024qwen2vlenhancingvisionlanguagemodels), and LongVU (wu2024longvideobench), which can process videos over an hour long. This highlights the growing need for effective benchmarks. In contrast, short video benchmarks have been widely explored, covering tasks like moment retrieval (liang2024tvrrankingdatasetrankedvideo; lei2021qvhighlightsdetectingmomentshighlights), action classification (caba2015activitynet; kinetics-400; soomro2012ucf101dataset101human), video reasoning (xie2024funqasurprisingvideocomprehension; xiao2021nextqanextphasequestionansweringexplaining), and video QA (xu2017video; lei2019tvqa; jang2017tgifqa). Among longer video benchmarks, TVQA (lei2019tvqa) is an early example with over 152.5 K QA pairs across 21.8 K clips (1.27 min avg.), totaling 461.2 hours. More recent datasets like MovieChat-1K(song2023moviechat) offer 1,000 movie clips (9.4 min avg.) with 14K annotations. CRAFT(ates2020craft) and GazeVQA (ilaslan2023gazevqa) extend to causal and embodied reasoning. Other long-form efforts include MLVU (zhou2024mlvucomprehensivebenchmarkmultitask), MoVQA (zhang2023movqa), and Video-MME (videomme), though their average durations peak at 17 minutes. LVBench(wang2024lvbench) stands out with 1-hour videos, but includes only 103 clips totaling 117 hours. Domain-specific datasets like Ego4D QA on EgoSchema(mangalam2023egoschema) focus on first-person views but use short (3-minute) clips. Motivated by these gaps, our benchmark evaluates grounding (temporal/visual anchoring) and reasoning (causality, narrative) skills to reflect human-like video understanding. As shown in Table 1, our dataset supports near-hour-long videos and offers the most significant total duration ($\sim$1K hours) and the largest QA-pair count in terms of long video understanding (87.7K).
@@ 3.2.2 Generating Question-Answer Pairs
Why Transcripts? Transcripts offer richer context than subtitles, including scene descriptions and character actions (e.g., “Monica and Chandler enter the café”). These cues allow for generating deeper, temporally grounded questions that reflect both visual and structural aspects of the story. For more details, see B.4 in the supplementary. For the TV shows, transcripts were sourced from reputable platforms such as foreverdreaming and manually verified and aligned with the video. For movies, we utilize the transcripts provided by MovieNet huang2020movienet, which have been pre-processed to ensure accurate alignment with both dialogue and subtitles, as detailed in Section A.5 huang2020movienet.
@@
Character Tracking: To track character actions in the video, we input the transcript and the video summary into GPT-4o, which generates question-answer pairs focused on identifying the sequence of actions performed by the main characters. Deep Context Understanding: For the deep context understanding question-answer pairs, we employed GPT-4o with the transcript and the summary to generate questions about the hidden motivations and social cues. Global Appearance: Our pipeline, as shown in Figure 2, integrates person detection khanam2024yolov11overviewkeyarchitectural and face-matching using InsightFace guo2021sample to identify characters in video frames. Reference images of the main cast were collected from IMDb and matched with detected faces, using an empirically determined threshold to filter out unmatched individuals. Cropped character images were saved chronologically in timestamped folders. To identify unique outfits, we extracted visual features using DINO-v2 oquab2024dinov2learningrobustvisual and removed redundant frames based on similarity scores, reducing approximately 200 frames to 20 while conservatively preserving visual diversity. The remaining images were then processed with GPT-4o GPT-4o to generate outfit descriptions. Multiple-choice questions were constructed by varying the sequence of outfits. In cases where a character wore a single outfit throughout, distractor options featuring incorrect outfits were introduced to maintain question complexity. Summarization & Spoiler Questions: We curate expert-written IMDb imdb summaries for human-level summarization. Spoiler questions are sourced from flagged plot discussions, focusing on twists or endings.
@@ 3.2.3 Data Integrity and Diversity
1) Redundancy Removal. To avoid question duplication, we compute pairwise textual similarity using BGE-M3 embeddings bge-m3 and remove highly similar questions (see Section B.5 for details). Additionally, since movies often contain over 100 events, many of which are redundant, we apply the same similarity-based filtering to retain only 20 diverse and representative events per movie.
@@ 3.3 Data Statistics
InfiniBench ultimately comprises 87.7K questions over videos averaging 53 minutes, with some running as long as 201 minutes, making it the largest and most challenging benchmark for long-video understanding. InfiniBench includes 923 episodes from six TV shows (lei2019tvqa) based on the long-form version in GoldFish ataallah2024goldfishvisionlanguageunderstandingarbitrarily and filtered 296 movies that have subtitles from MovieNet (huang2020movienet). These movies offer a broad and challenging foundation for benchmarking long-video understanding. Figures 4 and 4 illustrate the per-skill distribution, source breakdown (TVQA (lei2019tvqa), MovieNet (huang2020movienet)), and duration statistics across the dataset. We humanly verified 20% of the data, ensuring fair coverage across TV shows and movies. The remaining 80 % is available for training video models. The verified portion is further divided into test and validation splits, 10% each. All reported results and ablation studies are conducted using the test split.
@@ 4.1 Models
Open-Source Models We benchmark eleven leading open-source MLLMs, each evaluated using their official implementations and following their official best practices and configuration, e.g., recommending sampling strategy and frame-rate. These models vary in architecture, context length, and vision-language alignment strategies, offering a diverse perspective on the current research landscape: Qwen2.5-VL bai2025qwen25vltechnicalreport, Qwen2-VL wang2024qwen2vlenhancingvisionlanguagemodels, InternVL3 zhu2025internvl3, InternVL2.5 InternVL2.5, InternVL2InternVL2, LLaVA-OneVision (li2024llavaonevisioneasyvisualtask), LongVU (shen2024longvuspatiotemporaladaptivecompression),VideoChat-Flash li2025videochatflashhierarchicalcompressionlongcontext, InternLM-XComposer-2.5 zhang2024internlmxcomposer25versatilelargevision, Goldfish ataallah2024goldfishvisionlanguageunderstandingarbitrarily, and MiniGPT4-videominigpt4_video
Commercial Models We also evaluate two cutting-edge proprietary models via their official APIs: GPT-4o GPT-4o and Gemini 2.0 Flash  gemini_flash. These models represent the frontier of closed-access MLLMs with advanced capabilities in video, text, and multi-turn interaction. Their inclusion offers a valuable benchmark ceiling for current open-source efforts.
@@ 4.4.1 Self-enhancement Bias
The main results in Table 3 use GPT-4o as the judge model, which raises the potential concern of self-enhancement bias zheng2023judging, since GPT-4o serves as both the task model and evaluator. However, we argue that such bias is unlikely in our setup due to the clear separation between generation and evaluation. Specifically, during inference, the model accesses both video and subtitles to generate answers, whereas the judge model evaluates responses using only the generated text and reference answers. This separation in goals, modalities, and inputs minimizes the risk of any model-specific advantage transferring across tasks.
@@ 6 Limitations
Despite its strengths, our question–answer generation pipeline currently depends on the availability of transcripts, limiting its use to movies and TV shows. To broaden applicability, we will develop a more general pipeline that can handle arbitrary videos without preexisting transcripts. We also plan to leverage our dense, skill‐based annotations to create multi‐video benchmarks, such as season‐level evaluations, that extend each skill across related clips. Furthermore, our training split can support retrieval‐based MLLM methods that focus on relevant frames (inspired by Video-XL (liu2025video)), and we aim to explore long-video reasoning approaches, such as those based on the GRPO algorithm shao2024deepseekmathpushinglimitsmathematical, using samples from our dataset.
7 Acknowledgment
We would like to express our sincere gratitude to Habib Slim for his valuable assistance in preparing the visualizations accompanying this paper. His attention to detail and artistic insight significantly enhanced the overall quality and clarity of the work.
Models Filter Video Subtitles
Meta
data
Grounding Skills Reasoning Skills
Avg.
Acc.
Avg.
Score
Global
Appearance
Scene
Transitions
Character
Actions
Chronological
Understanding
Summar-
ization
Deep Context
Understanding
Spoiler
Understanding
Linking
Events
Random Per 20.0 20.0 20.0 20.0 N/A N/A N/A N/A 20.0 N/A
GPT4o w/o Filtering 51.8 42.9 43.5 77.2 6.3 6.8 6.7 7.5 53.9 6.8
29.1 39.6 46.8 82.0 7.2 7.0 5.9 7.5 49.4 6.9
25.7 26.1 39.1 56.5 3.4 5.2 3.9 7.4 36.8 5.0
w Filtering 49.7 37.7 39.9 61.0 6.3 6.4 6.6 6.8 47.1 6.5
26.5 34.2 42.2 68.5 7.1 6.5 5.8 6.8 42.8 6.5
22.2 19.9 31.9 35.4 3.3 4.8 3.5 6.6 27.4 4.6
Qwen2.5VL w/o Filtering 32.8 30.1 30.0 41.4 3.4 4.8 3.7 6.5 33.6 4.6
22.7 36.2 31.0 56.9 4.8 5.2 3.6 6.6 36.7 5.1
19.0 26.3 28.9 43.5 2.2 3.7 3.5 6.2 29.4 3.9
w Filtering 30.0 25.3 22.7 20.4 3.3 4.3 3.4 5.4 24.6 4.1
18.4 24.2 21.5 23.9 4.2 4.1 3.7 5.4 21.9 4.3
15.7 19.6 19.8 17.2 2.2 3.4 3.1 4.9 18.1 3.4
InternVL 3.0 w/o Filtering 35.9 27.8 24.8 41.2 3.9 4.1 3.6 6.2 32.4 4.4
30.8 34.3 24.8 46.5 4.2 4.2 3.8 6.4 34.0 4.7
25.0 25.6 26.5 38.9 1.9 3.6 3.9 6.1 29.0 3.9
w Filtering 34.3 27.8 20.5 31.1 3.8 3.7 3.3 5.2 28.4 4.0
29.2 31.2 20.8 33.9 4.2 3.9 3.5 5.4 28.8 4.2
23.4 23.3 23.3 24.9 2.0 3.3 3.6 5.0 23.7 3.5
LlaVa-OneVision w/o Filtering 23.2 27.6 27.0 41.1 2.0 4.0 3.2 6.1 29.7 3.9
20.8 27.8 29.8 43.4 4.0 5.1 3.7 6.6 30.4 4.8
16.7 27.8 28.5 39.3 0.8 3.7 3.4 5.8 28.1 3.4
w Filtering 21.1 24.7 23.3 30.3 2.0 3.7 2.7 5.1 24.8 3.4
18.5 24.5 25.7 28.9 3.9 4.5 3.3 5.6 24.4 4.3
14.8 24.5 26.2 28.4 0.8 3.4 2.9 4.8 23.5 3.0
Table 6: Analysis of the impact of filtering our benchmark by excluding the models’ internal knowledge and emphasizing the visual questions. Random Per. is random performance.
Models
Frame
Rate
Grounding Skills Reasoning Skills
Avg.
Acc.
Avg.
Score
Overall
Global
Appearance
Scene
Transitions
Character
Actions
Chronological
Understanding
Summar-
ization
Deep Context
Understanding
Spoiler
Understanding
Linking
Events
Random Per. N/A 19.48 16.67 16.52 40.56 N/A N/A N/A N/A 23.31 N/A N/A
GPT-4o 450 FPV 54.26 50.0 43.96 49.66 6.2 6.49 5.61 7.00 49.47 6.33 56.36
Gemini-2.0 1 FPS 50.39 50.79 56.61 36.60 5.70 5.40 3.41 5.94 48.60 5.11 49.86
InternVL3 128 FPV 33.33 28.57 23.46 36.29 3.91 3.79 3.41 5.28 30.41 4.10 35.69
Qwen2VL 768 FPV 31.01 23.81 32.59 48.41 2.26 4.23 2.93 5.14 33.96 3.64 35.18
Qwen2.5VL 768 FPV 37.98 22.22 22.53 27.17 3.22 4.13 3.67 5.34 27.48 4.09 34.19
LLava-OneVision 128 FPV 37.21 25.40 20.11 43.91 1.99 3.66 3.28 5.21 31.66 3.54 33.50
Goldfish 60 FPW 20.93 12.70 21.97 38.48 2.88 4.85 3.61 5.48 23.52 4.21 32.79
InternVL2 128 FPV 24.03 15.87 22.16 42.16 2.74 3.37 2.87 5.02 26.06 3.50 30.53
VideoChat-Flash 1000 FPV 28.68 20.63 33.33 29.73 2.45 3.24 2.35 4.20 28.09 3.06 29.35
InternVL2.5 128 FPV 34.88 20.63 21.6 32.79 2.60 2.78 2.35 4.20 27.48 2.98 28.65
InternLM-XComposer 16 FPW 32.56 14.29 28.49 30.11 1.72 2.50 2.11 4.13 26.36 2.62 26.26
MiniGPT4-video 60 FPV 17.83 11.11 27.00 40.1 2.04 2.87 2.15 3.39 24.01 2.61 25.07
LongVU 512 FPV 20.93 15.87 21.42 29.86 1.70 2.84 3.07 3.53 22.02 2.79 24.94
Table 7: InfiniBench leaderboard on the validation-set, across eight skills. Models are ranked by their equally weighted performance for Grounding and reasoning skills using this formula $\left(0.5\cdot\frac{\text{acc}}{100}+0.5\cdot\frac{\text{score}}{10}\right)$. FPV (Frames Per Video), FPS (Frames Per Second), and FPW (Frames Per Window) are reported. All models use both video and subtitles as inputs. Random Per. is random performance.
Models
Frame
Rate
Grounding Skills Reasoning Skills
Avg.
Acc.
Avg.
Score
Overall
Global
Appearance
Scene
Transitions
Character
Actions
Chronological
Understanding
Summar-
ization
Deep Context
Understanding
Spoiler
Understanding
Linking
Events
Random Per. N/A 20.00 20.00 20.00 20.00 N/A N/A N/A N/A 20.00 N/A N/A
GPT-4o 450 FPV 51.83 42.91 43.49 77.15 6.32 6.82 6.73 7.47 53.85 6.84 61.10
Gemini-2.0 1 FPS 46.73 44.59 55.94 67.62 5.80 6.56 4.36 6.19 53.72 5.73 55.50
Qwen2.5VL 768 FPV 32.78 30.10 30.04 41.36 3.35 4.84 3.71 6.46 33.57 4.59 39.74
InternVL3 128 FPV 35.93 27.76 24.75 41.17 3.86 4.06 3.57 6.19 32.40 4.42 38.30
Qwen2VL 768 FPV 25.37 31.37 34.91 36.68 2.23 4.87 3.59 6.03 32.08 4.18 36.94
Goldfish 60 FPW 17.28 24.45 24.03 29.32 3.02 5.45 3.61 6.56 23.77 4.66 35.19
VideoChat-Flash 1000 FPV 21.48 31.37 37.48 48.14 2.64 3.93 2.38 5.10 34.62 3.51 34.87
LLava-OneVision 128 FPV 23.15 27.57 27.04 41.07 2.04 4.04 3.18 6.12 29.71 3.85 34.08
InternVL2 128 FPV 27.72 26.18 23.89 30.47 2.89 3.74 3.16 5.96 27.07 3.94 33.22
InternVL2.5 128 FPV 28.58 27.10 25.18 34.00 2.46 3.10 2.30 5.10 28.72 3.24 30.56
InternLM-XComposer 16 FPW 22.53 30.33 30.33 34.48 1.63 2.86 2.43 5.02 29.42 2.99 29.63
LongVU 512 FPV 27.04 22.38 25.89 21.49 1.70 3.26 3.00 4.19 24.20 3.04 27.29
MiniGPT4-video 60 FPV 18.52 25.84 29.18 25.79 2.09 3.09 2.21 3.89 24.83 2.82 26.52
Table 8: InfiniBench before filtration, Section 3.2.3, leaderboard across eight skills. Models are ranked by their equally weighted performance for Grounding and reasoning skills using this formula $\left(0.5\cdot\frac{\text{acc}}{100}+0.5\cdot\frac{\text{score}}{10}\right)$. FPV (Frames Per Video), FPS (Frames Per Second), and FPW (Frames Per Window) are reported. All models use both video and subtitles as inputs. Random Per. is random performance.
Models # Frames
I don’t
know option
Grounding Skills
Global
Appearance
Scene
Transitions
Character
Actions
Chronological
Understanding
Avg.
Acc.
Qwen2.5-VL 768 30.0 25.3 22.7 20.3 24.6
31.4 23.3 23.9 24.3 25.7
InternVL3.0 128 34.3 27.8 20.5 31.1 28.4
35.5 30.9 22.9 33.5 30.7
Qwen2-VL 768 23.5 28.2 30.2 27.4 27.3
25.2 31.8 29.9 32.3 29.8
Video-flash 1000 20.5 29.6 34.9 37.4 30.6
20.8 27.4 31.9 36.8 29.2
InternLM-XC 16 FPW 20.1 27.9 26.6 26.4 25.3
22.6 30.1 25.5 26.6 26.2
InternVL2.5 128 27.1 25.1 21.2 29.2 25.6
27.7 27.7 22.2 28.9 26.7
LongVU 512 27.7 21.0 28.0 19.0 23.9
27.7 23.5 29.5 23.5 26.0
Table 9: Analysis of the impact of incorporating the “I don’t know” option on model performance.